OpenStudy (cooltowl):

Verify the identity cotx-tanx all over cot^2x- tan^2x= sinxcosx

OpenStudy (anonymous):

How would you go to factor ( cot^2x- tan^2x)? Factoring this will make your life much easier.

OpenStudy (anonymous):

(cotx-tanx)/(cot^2x- tan^2x) =(cotx-tanx)/ ( (cotx-tanx) * (cotx+tanx) ) =1/(cotx+tanx)

OpenStudy (anonymous):

1/( (cosx/sinx) + (sinx/cosx) ) =1/ ( (cos^2x+ sin^2x) / (cosx*sinx) ) = cosx*sinx/(cos^2x+ sin^2x) = cosxsinx as we know (cos^2x+ sin^2x)=1